# references.bib

### v1

@@ 3.2 Convergence of minimizers
Of particular interest to us is the case $\lambda=0$ for which minimizers of $\mathcal{L}^{0}_{f}$ are related to the teacher measure $\bar{\nu}$ by eq. 17. However in practice, minimization of $\mathcal{L}^{\lambda}_{f}$ is easier in the presence of a regularization parameter $\lambda>0$. For this reason, we are interested in the asymptotic behavior of minimizers of $\mathcal{L}^{\lambda}_{f}$ when $\lambda\to 0^{+}$.
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\xrightarrow{k\to\infty}\mathcal{L}^{0}_{f}(\mu).$
Then by the dual formulation of $\mathcal{L}^{\lambda}_{f}$ in eq. 19, for every $n,k\geq 0$:
$\mathcal{L}^{\lambda_{n}}(\mu_{n})$ $\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu_{n}+\langle\alpha_{k},Y\rangle-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}$
$=-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}(\mu_{n}-\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\rangle$
$\geq-\left\|f^{*}(\Phi^{\top}\alpha_{k})\right\|_{\mathcal{C}^{0,1}}\mathcal{W}_{1}(\mu_{n},\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\langle\alpha_{k},Y\rangle\rangle.$
But then, since $\mathcal{W}_{1}(\mu_{n},\mu)\to 0$ and $\lambda_{n}\to 0$, one can find a sequence $(k_{n})_{n\geq 0}$ diverging to infinity such that it holds:
@@
Thus, for every $n\geq 0$, $\mathcal{L}^{\lambda_{n}}_{f}(\mu_{n})\geq\mathcal{L}^{0}_{f}(\mu)+o(1)$ and the result follows. ∎
@@ 4 Training with gradient flow
In the rest of this work we consider the optimization over the feature distribution $\mu\in\mathcal{P}(\Omega)$ for the minimization of the reduced risk $\mathcal{L}^{\lambda}_{f}$, for $\lambda\geq 0$. Specifically, we consider a gradient flow algorithm. In the case of a finite number of feature $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ such gradient flow is defined as the solution of the equation:
@@
$\partial_{t}\mu_{t}-\mathrm{div}\left(\mu_{t}\nabla\frac{\delta\mathcal{L}^{\lambda}_{f}}{\delta\mu}[\mu_{t}]\right),\quad\text{over $(0,\infty)\times\Omega$,}$ (29)
@@
We study in this section the well-posedness of the above Wasserstein gradient flow equation by distinguishing the case where $\lambda>0$ and the case $\lambda=0$. In the latter case, we show the Wasserstein gradient flow corresponds to an ultra-fast diffusion equation.
@@ 4.1 Wasserstein gradient flows in the case $\lambda>0$
with $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ the maximizer in eq. 19. We show that the gradient field of this potential indeed defines a notion of “gradient” for the functional $\mathcal{L}^{\lambda}_{f}$ w.r.t. the Wasserstein-$2$ topology on $\mathcal{P}(\Omega)$. More precisely, absolutely continuous curves $(\mu_{t})_{t\in[0,1]}$ in the space $\mathcal{P}(\Omega)$ are characterised as solutions to a continuity equation:
$\partial_{t}\mu_{t}+\mathrm{div}(\mu_{t}v_{t})=0\quad\text{over $(0,1)\times\Omega$}$ (31)
for some velocity field $v$ such that $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}((0,1))$ [santambrogio2015optimal, Thm. 5.14]. This equation has to be understood in the sense of distributions, that is in duality with the set $\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega)$ of compactly supported functions, i.e.:
$\int_{0}^{1}\int_{\Omega}\left(\partial_{t}\varphi+\langle\nabla\varphi,v_{t}\rangle\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,1)\times\Omega).$ (32)
@@
Assume $\phi\in L^{2}(\rho,\mathcal{C}^{1})$, $f$ satisfies Assumption 2 with $f^{*}\in\mathcal{C}^{1}_{\mathrm{loc}}(\mathbb{R})$ and consider $\lambda>0$. Let $(\mu_{t})_{t\in[0,1]}$ be an absolutely continuous curve in $\mathcal{P}(\Omega)$ solution of the continuity equation eq. 32 for some velocity field $v$ such that $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}((0,1))$. Then $(\mathcal{L}^{\lambda}_{f}(\mu_{t}))_{t\in[0,1]}$ is absolutely continuous and for a.e. $t^{\prime},t\in[0,1]$:
$\mathcal{L}^{\lambda}_{f}(\mu_{t^{\prime}})-\mathcal{L}^{\lambda}_{f}(\mu_{t})=\int_{t}^{t^{\prime}}\langle\nabla\mathcal{L}^{\lambda}_{f}[\mu_{s}],v_{s}\rangle_{L^{2}(\mu_{s})}\mathrm{d}s$
where, for $\mu\in\mathcal{P}_{2}(\Omega)$, the velocity field $\nabla\mathcal{L}^{\lambda}_{f}[\mu]\in L^{2}(\mu)$ is defined as:
$\forall\omega\in\Omega,\quad\nabla\mathcal{L}^{\lambda}_{f}[\mu](\omega)\coloneqq-\nabla\left(f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu])\right)(\omega)$
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}\mathcal{V}_{\alpha}(\mu)-\frac{\lambda}{2}\|\alpha\|^{2}+\langle\alpha,Y\rangle,$
@@
In particular, at fixed $\alpha\in L^{2}(\rho)$, it follows from the assumptions on $\phi$ and $f^{*}$ that the potential $f^{*}(\Phi^{\top}\alpha)$ is in $\mathcal{C}^{1}(\Omega)$ with $\|f^{*}(\Phi^{\top}\alpha)\|_{\mathcal{C}^{1}}\leq C(\|\alpha\|_{L^{2}(\rho)})$ for some continuous function $C$. Thus, by properties of the continuity equation, $\mathcal{V}_{\alpha}(\mu_{t})$ is absolutely continuous and its derivative is give for a.e. $t\in[0,1]$ by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{V}_{\alpha}(\mu_{t})=-\int_{\Omega}\langle\nabla f^{*}(\Phi^{\top}\alpha),v_{t}\rangle\mathrm{d}\mu_{t}.$
@@
$\partial_{t}\mu_{t}-\mathrm{div}(\mu_{t}\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}])=0\quad\text{over $(0,\infty)\times\Omega$}.$ (33)
@@
Definition 4.1 (Gradient flow of $\mathcal{L}^{\lambda}_{f}$).
Let $\mu_{0}\in\mathcal{P}_{2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting at $\mu_{0}$ if it is a locally absolutely continuous curve s.t. $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$ and satisfying the continuity equation eq. 33 in the sense of distribution, i.e.:
$\int_{0}^{\infty}\int_{\Omega}\left(\partial_{t}\varphi-\nabla\varphi\cdot\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega).$ (34)
@@
Note that, in the case where $\Omega$ is a closed, bounded, smooth and convex domain, our definition eq. 32 of solutions to the continuity equation enforces no-flux conditions on the boundary $\partial\Omega$. Indeed we consider test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,1)\times\Omega)$ that can be supported on the whole domain $\Omega$ (which is always assumed closed). Thus, eq. 32 enforces $\langle\mu_{t}v_{t},\vec{n}\rangle=0$ in the sense of distribution, where $\vec{n}$ is the outer normal vector to the boundary $\partial\Omega$.
In case of the gradient flow equation eq. 34, this boundary condition is for example satisfied if one assumes $\langle\nabla_{\omega}\phi(\omega,x),\vec{n}\rangle=0$ for every $x\in\mathbb{R}^{d}$ and every $\omega\in\partial\Omega$. Another way of ensuring the no-flux condition is to remove the outer part of the gradient field $\nabla\mathcal{L}^{\lambda}_{f}$ on the boundary $\partial\Omega$, which can be performed by clipping the features.
@@
To show the well-posedness of gradient flows, we rely on convexity properties of the functional $\mathcal{L}^{\lambda}_{f}$. Indeed, by the dual formulation in eq. 19, we can express $\mathcal{L}^{\lambda}_{f}$ as a supremum over semi-convex functional. As a consequence, the Lemma 4.2 below shows that, for $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ is semi-convex along (generalized) geodesics of the Wasserstein space (see [ambrosio2008gradient, Def. 9.2.4] for the definiton of generalized geodesics). However, note that such an argument can not be extended to the case $\lambda=0$ since the semi-convexity constant blows-up when $\lambda\to 0^{+}$. For example, in the case $f(t)=|t|^{2}$ this constant scales as $\lambda^{-2}$.
Lemma 4.2 (Geodesic semi-convexity).
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)(\omega)\mathrm{d}\mu(\omega)-\frac{\lambda}{2}\|\alpha\|^{2}+\langle\alpha,Y\rangle,$
@@
Then, from the assumptions on $f^{*}$, the composition $f^{*}(\Phi^{\top}\alpha)$ is also in $\mathcal{C}^{1,1}(\Omega)$ and by [ambrosio2008gradient, Prop.9.3.2] the functional $\mu\mapsto\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu$ is $C$-semiconvex along generalized geodesic for some constant $C=C(f^{*},\|\alpha\|_{L^{2}(\rho)}\|\phi\|_{L^{2}(\rho,\mathcal{C}^{1,1})})$. Moreover, similarly as in the proof of Lemma 4.1, one can restrict the definition of $\mathcal{L}^{\lambda}_{f}$ to the supremum over $\alpha\in L^{2}(\rho)$ with $\|\alpha\|_{L^{2}(\rho)}\leq R_{\lambda}$. The result then follows by taking a supremum over (uniformly) semi-convex functionals. ∎
The semi-convexity of $\mathcal{L}^{\lambda}_{f}$ along generalized geodesic ensures the existence and uniqueness of gradient flows in the sense of Definition 4.1.
Theorem 4.1 (Well-posedness of the gradient flow equation for $\lambda>0$).
Assume the assumptions of Lemma 3.1 hold. Then for any $\lambda>0$ and any initialization $\mu_{0}\in\mathcal{P}_{2}(\Omega)$ there exists a unique gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ in the sense of Definition 4.1. Moreover, if $(\mu_{t})_{t\geq 0},(\mu_{t}^{\prime})_{t\geq 0}$ are gradient flows for $\mathcal{L}^{\lambda}_{f}$ with respective initializations $\mu_{0},\mu_{0}^{\prime}\in\mathcal{P}(\Omega)$ then for every $t\geq 0$:
@@
The chain rule formula established in Lemma 4.1 shows that for every $\mu\in\mathcal{P}(\Omega)$ the vector field $\mathcal{L}^{\lambda}_{f}[\mu]$ is a strong subdifferential of $\mathcal{L}^{\lambda}_{f}$ in the sense of [ambrosio2008gradient, Def. 10.3.1 and eq. (10.3.12)]. Existence, uniqueness and properties of the gradient flows then follows from the geodesic semi-convexity of $\mathcal{L}^{\lambda}_{f}$ established in Lemma 4.2 and the application of [ambrosio2008gradient, Def. 11.2.1]
Finally, it is a classical property of weak solutions to continuity equations that the gradient flows of $\mathcal{L}^{\lambda}_{f}$ can be represented in terms of push-forward by a flow map.
@@
$\forall t\geq 0,\quad\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\circ X_{t},\quad\text{with}\quad X_{0}=\mathrm{Id}_{\Omega}.$
In particular, if $(\omega_{i}(t))_{t\geq 0}$ for $i\in\{1,...,M\}$ are solutions to eq. 28 then the empirical distribution $\hat{\mu}_{t}\coloneqq\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}(t)}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ in the sense of Definition 4.1 and thus $\omega_{i}(t)=X_{t}(\omega_{i}(0))$ for $i\in\{1,...,M\}$ and $t\geq 0$.
@@
For every $t\geq 0$, similarly as in the proof of Lemma 4.1 we have that the dual variable is bounded by $\|\alpha^{\lambda}_{f}[\mu_{t}]\|_{L^{2}(\rho)}\leq R_{\lambda}$ and from the assumption on the regularity of $\phi$ it follows that:
$\|f^{*}(\Phi^{\top}\alpha)\|_{\mathcal{C}^{1,1}}\leq C.$
for some constant $C=C(f^{*},\lambda)$. Then by definition $\nabla\mathcal{L}^{\lambda}[\mu_{t}]=-\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu_{t}])\in\mathcal{C}^{0,1}$ and the first part of the result follows from classical results of ODE theory [hale2009ordinary] and on representation of solutions to continuity equation [ambrosio2008gradient, Thm. 8.1.8]. For the second part of the result, it suffices to remark that by the definition of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ and $\nabla\mathcal{L}^{\lambda}_{f}$, for $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and $j\in\{1,...,M\}$:
@@
where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. Therefore by eq. 28 we have that for a test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,\infty)\times\Omega)$:
@@
meaning $(\hat{\mu}_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ according to Definition 4.1. ∎
@@
In the case where $\lambda>0$, associating the contraction rate of the gradient flow obtained in Theorem 4.1 with classical result on the approximation of measure by empirical distribution we obtain an approximation result for the minimization of $\mathcal{L}^{\lambda}_{f}$ with a finite number of features. For conciseness, we only state the result in the case $d\geq 3$, but similar results hold for $d\in\{1,2\}$.

### v2

@@ 3.2 Convergence of minimizers
Of particular interest to us is the case $\lambda=0$ for which minimizers of $\mathcal{L}^{0}_{f}$ are related to the teacher measure $\bar{\nu}$ by Eq. 25. However, in practice, minimization of $\mathcal{L}^{\lambda}_{f}$ is easier in the presence of a regularization parameter $\lambda>0$. For this reason, we are interested in the asymptotic behavior of minimizers of $\mathcal{L}^{\lambda}_{f}$ when $\lambda\to 0^{+}$.
@@
$-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\left<\alpha_{k},Y\right>\xrightarrow{k\to\infty}\mathcal{L}^{0}_{f}(\mu).$
Then, by the dual formulation of $\mathcal{L}^{\lambda}_{f}$ in Eq. 27, for every $n,k\geq 0$:
$\mathcal{L}^{\lambda_{n}}(\mu_{n})$ $\geq-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu_{n}+\left<\alpha_{k},Y\right>-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}$
$=-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}(\mu_{n}-\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\left<\alpha_{k},Y\right>$
$\geq-\left\|f^{*}(\Phi^{\top}\alpha_{k})\right\|_{\mathcal{C}^{0,1}}\mathcal{W}_{1}(\mu_{n},\mu)-\frac{\lambda_{n}}{2}\|\alpha_{k}\|^{2}_{L^{2}{\rho}}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha_{k})\mathrm{d}\mu+\left<\alpha_{k},Y\right>.$
But then, since $\mathcal{W}_{1}(\mu_{n},\mu)\to 0$ and $\lambda_{n}\to 0$, one can find an increasing sequence $(k_{n})_{n\geq 0}$ s.t.:
@@
Thus $\mathcal{L}^{\lambda_{n}}_{f}(\mu_{n})\geq\mathcal{L}^{0}_{f}(\mu)+o(1)$ for every $n\geq 0$ and the result follows. ∎
@@ 4 Training with gradient flow
In the rest of this work we consider the optimization over the feature distribution $\mu\in\mathcal{P}(\Omega)$ for the minimization of the reduced risk $\mathcal{L}^{\lambda}_{f}$, for $\lambda\geq 0$. Specifically, we consider a gradient flow algorithm. In the case of a finite number of features $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ such gradient flow is defined as the solution of the equation:
@@
$\partial_{t}\mu_{t}-\mathrm{div}\left(\mu_{t}\nabla\frac{\delta\mathcal{L}^{\lambda}_{f}}{\delta\mu}[\mu_{t}]\right)=0,\quad\text{on $(0,\infty)\times\Omega$,}$ (37)
@@
We study in this section the well-posedness of the above Wasserstein gradient flow equation by distinguishing the case where $\lambda>0$ and the case $\lambda=0$. In the latter case, we show the Wasserstein gradient flow corresponds to a weighted ultra-fast diffusion equation [iacobelli2019weighted].
@@ 4.1 Wasserstein gradient flows in the case $\lambda>0$
with $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ the maximizer in Eq. 27. We show that the gradient field of this potential indeed defines a notion of “gradient” for the functional $\mathcal{L}^{\lambda}_{f}$ w.r.t. the Wasserstein topology on $\mathcal{P}(\Omega)$.
Locally absolutely continuous curves $(\mu_{t})_{t\in(0,+\infty)}$ in the space $\mathcal{P}(\Omega)$, equipped with the Wasserstein distance $\mathcal{W}_{2}$, are characterised as solutions to the continuity equation:
$\partial_{t}\mu_{t}+\mathrm{div}(\mu_{t}v_{t})=0\quad\text{on $(0,+\infty)\times\Omega$}$ (39)
for some velocity field $v$ such that $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}_{\mathrm{loc}}((0,+\infty))$ [santambrogio2015optimal, Thm. 5.14]. This equation has to be understood in the sense of distributions, that is in duality with the set $\mathcal{C}^{\infty}_{c}((0,+\infty)\times\Omega)$ of compactly supported test functions, i.e.:
$\int_{0}^{1}\int_{\Omega}\left(\partial_{t}\varphi+\left<\nabla\varphi,v_{t}\right>\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,+\infty)\times\Omega).$ (40)
@@
Assume $\phi\in L^{2}(\rho,\mathcal{C}^{1})$, $f$ satisfies Assumption 2 with $f^{*}\in\mathcal{C}^{1}_{\mathrm{loc}}(\mathbb{R})$ and consider $\lambda>0$. Let $(\mu_{t})_{t\in(0,+\infty)}$ be a locally absolutely continuous curve in $\mathcal{P}(\Omega)$ solution of the continuity equation Eq. 40 for a velocity field $v$ s.t. $\|v_{t}\|_{L^{2}(\mu_{t})}\in L^{1}_{\mathrm{loc}}((0,+\infty))$. Then $(\mathcal{L}^{\lambda}_{f}(\mu_{t}))_{t\in(0,+\infty)}$ is locally absolutely continuous and for a.e. $t^{\prime},t\in(0,+\infty)$:
$\mathcal{L}^{\lambda}_{f}(\mu_{t^{\prime}})-\mathcal{L}^{\lambda}_{f}(\mu_{t})=\int_{t}^{t^{\prime}}\left<\nabla\mathcal{L}^{\lambda}_{f}[\mu_{s}],v_{s}\right>_{L^{2}(\mu_{s})}\mathrm{d}s\,,$
where for $\mu\in\mathcal{P}_{2}(\Omega)$ the velocity field $\nabla\mathcal{L}^{\lambda}_{f}[\mu]\in L^{2}(\mu)$ is defined by:
$\forall\omega\in\Omega,\quad\nabla\mathcal{L}^{\lambda}_{f}[\mu](\omega)\coloneqq-\nabla\left(f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu])\,,\right)(\omega)$
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}\mathcal{V}_{\alpha}(\mu)-\frac{\lambda}{2}\|\alpha\|^{2}+\left<\alpha,Y\right>,$
@@
In particular, at fixed $\alpha\in L^{2}(\rho)$, it follows from the assumptions on $\phi$ and $f^{*}$ that the potential $f^{*}(\Phi^{\top}\alpha)$ is in $\mathcal{C}^{1}(\Omega)$ with $\|f^{*}(\Phi^{\top}\alpha)\|_{\mathcal{C}^{1}}\leq C(\|\alpha\|_{L^{2}(\rho)})$ for some continuous function $C$. Thus, by properties of the continuity equation, $\mathcal{V}_{\alpha}(\mu_{t})$ is locally absolutely continuous and its derivative is given for a.e. $t\in(0,+\infty)$ by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{V}_{\alpha}(\mu_{t})=-\int_{\Omega}\left<\nabla f^{*}(\Phi^{\top}\alpha),v_{t}\right>\mathrm{d}\mu_{t}.$
@@
$\partial_{t}\mu_{t}-\mathrm{div}(\mu_{t}\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}])=0\quad\text{on $(0,+\infty)\times\Omega$}.$ (41)
@@
Definition 1 (Gradient flow of $\mathcal{L}^{\lambda}_{f}$).
Let $\mu_{0}\in\mathcal{P}_{2}(\Omega)$. We say $(\mu_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ if it is a locally absolutely continuous curve on $(0,+\infty)$ s.t. $\lim_{t\to 0^{+}}\mu_{t}=\mu_{0}$ and if it satisfies the continuity equation Eq. 41 in the sense of distribution, i.e.:
$\int_{0}^{\infty}\int_{\Omega}\left(\partial_{t}\varphi-\nabla\varphi\cdot\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\right)\mathrm{d}\mu_{t}\mathrm{d}t=0,\quad\forall\varphi\in\mathcal{C}^{\infty}_{c}((0,+\infty)\times\Omega).$ (42)
@@
Note that, in the case where $\Omega$ is a closed, bounded, smooth and convex domain, our definition Eq. 40 of solutions to the continuity equation enforces no-flux conditions on the boundary $\partial\Omega$. Indeed we consider test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,+\infty)\times\Omega)$ that can be supported on the whole domain $\Omega$ (which is always assumed closed). Thus, Eq. 40 enforces $\mu_{t}v_{t},\vec{n}\right>=0$ in the sense of distribution, where $\vec{n}$ is the outer normal vector to the boundary $\partial\Omega$.
In case of the gradient flow equation Eq. 42, this boundary condition is for example satisfied if one assumes $\nabla_{\omega}\phi(\omega,x),\vec{n}\right>=0$ for every $x\in\mathbb{R}^{d}$ and every $\omega\in\partial\Omega$. Another way of ensuring the no-flux condition is to remove the outer part of the gradient field $\nabla\mathcal{L}^{\lambda}_{f}$ on the boundary $\partial\Omega$, which can be performed by clipping the features.
@@
To show the well-posedness of gradient flows, we rely on convexity properties of the functional $\mathcal{L}^{\lambda}_{f}$. Indeed, by the dual formulation in Eq. 27, we can express $\mathcal{L}^{\lambda}_{f}$ as a supremum over semiconvex functionals. As a consequence, the Lemma 4.2 below shows that, for $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ is semiconvex along (generalized) geodesics of the Wasserstein space (see [ambrosio2008gradient, Def. 9.2.4] for the definition of generalized geodesics). However, note that such an argument can not be extended to the case $\lambda=0$ since the semiconvexity constant blows-up when $\lambda\to 0^{+}$. For example, in the case $f(t)=|t|^{2}$ this constant scales as $\lambda^{-2}$.
Lemma 4.2 (Geodesic semiconvexity).
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)(\omega)\mathrm{d}\mu(\omega)-\frac{\lambda}{2}\|\alpha\|^{2}+\left<\alpha,Y\right>,$
@@
Then, from the assumptions on $f^{*}$, the composition $f^{*}(\Phi^{\top}\alpha)$ is also in $\mathcal{C}^{1,1}(\Omega)$ and by [ambrosio2008gradient, Prop.9.3.2] the functional $\mu\mapsto\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu$ is $C$-semiconvex along generalized geodesics for some constant $C=C(f^{*},\|\alpha\|_{L^{2}(\rho)}\|\phi\|_{L^{2}(\rho,\mathcal{C}^{1,1})})$. Moreover, similarly as in the proof of Lemma 4.1, one can restrict the definition of $\mathcal{L}^{\lambda}_{f}$ to the supremum over $\alpha\in L^{2}(\rho)$ with $\|\alpha\|_{L^{2}(\rho)}\leq R_{\lambda}$. The result then follows by taking a supremum over (uniformly) semiconvex functionals. ∎
The semiconvexity of $\mathcal{L}^{\lambda}_{f}$ along generalized geodesics ensures the existence and uniqueness of gradient flows in the sense of Definition 1.
Theorem 1 (Well-posedness of the gradient flow equation for $\lambda>0$).
Assume the assumptions of Lemma 4.2 hold. Then for any $\lambda>0$ and any initialization $\mu_{0}\in\mathcal{P}_{2}(\Omega)$ there exists a unique gradient flow for $\mathcal{L}^{\lambda}_{f}$ starting from $\mu_{0}$ in the sense of Definition 1. Moreover, if $(\mu_{t})_{t\geq 0},(\mu_{t}^{\prime})_{t\geq 0}$ are gradient flows for $\mathcal{L}^{\lambda}_{f}$ with respective initializations $\mu_{0},\mu_{0}^{\prime}\in\mathcal{P}(\Omega)$ then for every $t\geq 0$:
@@
The chain rule formula established in Lemma 4.1 shows that for every $\mu\in\mathcal{P}(\Omega)$ the vector field $\mathcal{L}^{\lambda}_{f}[\mu]$ is a strong subdifferential of $\mathcal{L}^{\lambda}_{f}$ in the sense of [ambrosio2008gradient, Def. 10.3.1 and eq. (10.3.12)]. Existence, uniqueness and contractivity properties of the gradient flow then follow from the geodesic semiconvexity of $\mathcal{L}^{\lambda}_{f}$ established in Lemma 4.2 and the application of [ambrosio2008gradient, Def. 11.2.1]
Finally, it is a classical property of weak solutions to continuity equations that gradient flows of $\mathcal{L}^{\lambda}_{f}$ can be represented in terms of push-forward by a flow map.
@@
$\forall t\geq 0,\quad\frac{\mathrm{d}}{\mathrm{d}t}X_{t}=-\nabla\mathcal{L}^{\lambda}_{f}[\mu_{t}]\circ X_{t},\quad\text{with}\quad X_{0}=\mathrm{Id}_{\Omega}.$
In particular, if $(\omega_{i}(t))_{t\geq 0}$ for $i\in\{1,...,M\}$ are solutions to Eq. 36 then the empirical distribution $\hat{\mu}_{t}\coloneqq\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}(t)}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ in the sense of Definition 1 and thus $\omega_{i}(t)=X_{t}(\omega_{i}(0))$ for $i\in\{1,...,M\}$ and $t\geq 0$.
@@
For every $t\geq 0$, similarly as in the proof of Lemma 4.1, we have that the dual variable is bounded by $\|\alpha^{\lambda}_{f}[\mu_{t}]\|_{L^{2}(\rho)}\leq R_{\lambda}$ and from the assumption on the regularity of $\phi$ it follows that:
$\|f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu_{t}])\|_{\mathcal{C}^{1,1}}\leq C.$
for some constant $C=C(f^{*},\lambda)$. Then by definition $\nabla\mathcal{L}^{\lambda}[\mu_{t}]=-\nabla f^{*}(\Phi^{\top}\alpha^{\lambda}_{f}[\mu_{t}])\in\mathcal{C}^{0,1}$ and the first part of the result follows from classical results of ODE theory [hale2009ordinary] and on representation of solutions to continuity equations [ambrosio2008gradient, Thm. 8.1.8]. For the second part of the result it suffices to remark that, by the definition of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ and $\nabla\mathcal{L}^{\lambda}_{f}$, for $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and $j\in\{1,...,M\}$:
@@
where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. Therefore, by Eq. 36 we have for any test function $\varphi\in\mathcal{C}^{\infty}_{c}((0,+\infty)\times\Omega)$:
@@
meaning $(\hat{\mu}_{t})_{t\geq 0}$ is a gradient flow for $\mathcal{L}^{\lambda}_{f}$ according to Definition 1. ∎
@@
In the case where $\lambda>0$, associating the contraction rate of the gradient flow obtained in Theorem 1 with classical results on the approximation of measure by empirical distributions we obtain an approximation result for the minimization of $\mathcal{L}^{\lambda}_{f}$ with a finite number of features. For conciseness, we only state the result in the case $d\geq 3$, but similar results hold for $d\in\{1,2\}$.